What is an equation that has variables on both sides and have a solution of -3
Answer by Edwin McCravy(20056)


Start with the final answer you would want to
get, which is

    x = -3

and work backwards.

Make up something arbitrary to add to both sides
that has a variable and a constant to both sides,
like maybe 7x + 2.  Add that to both sides

    x + 7x + 2 = -3 + 7x + 2

That will do since it has variables on both sides.  

But you might want to collect like terms on both sides:

    8x + 2 = 7x - 1

You can make them up by the hundreds.

Edwin